# EVOLvE: Evaluating and Optimizing LLMs For Exploration

### v1

@@ Abstract
Despite their success in many domains, large language models (LLMs) remain under-studied in scenarios requiring optimal decision-making under uncertainty. This is crucial as many real-world applications, ranging from personalized recommendations to healthcare interventions, demand that LLMs not only predict but also actively learn to make optimal decisions through exploration. In this work, we measure LLMs’ (in)ability to make optimal decisions in bandits, a state-less reinforcement learning setting relevant to many applications. We develop a comprehensive suite of environments, including both context-free and contextual bandits with varying task difficulties, to benchmark LLMs’ performance. Motivated by the existence of optimal exploration algorithms, we propose efficient ways to integrate this algorithmic knowledge into LLMs: by providing explicit algorithm-guided support during inference; and through algorithm distillation via in-context demonstrations and fine-tuning, using synthetic data generated from these algorithms. Impressively, these techniques allow us to achieve superior exploration performance with smaller models, surpassing larger models on various tasks. We conducted an extensive ablation study to shed light on various factors, such as task difficulty and data representation, that influence the efficiency of LLM exploration. Additionally, we conduct a rigorous analysis of the LLM’s exploration efficiency using the concept of regret, linking its ability to explore to the model size and underlying algorithm.
@@ 1 Introduction
We study LLMs’ in-context exploration capabilities under the simplified framework of bandits — a stateless form of reinforcement learning that is highly applicable to many domains. We set up the LLM to interact with the environment over $T$ rounds. In each round, it receives the full history of its past interactions, the current state (if provided), and a set of actions, and it is tasked with selecting an action to maximize the cumulative reward. Ideally, the LLM should adaptively choose an action in each round to learn the reward distributions of different actions and eventually converge to consistently selecting the optimal one. We study LLM’s ability to do so in-context, without the need to re-train, which we dubbed as in-context exploration [^1] Unlike random or reward-free exploration, in-context exploration represents a form of in-context self-improvement, where the LLM gradually enhances its ability to solve a task through effective in-context exploration..
We introduce BanditBench, a comprehensive suite of multi-armed bandit (MAB) (Slivkins et al., 2019) and contextual bandit (CB) (Li et al., 2010) environments in natural language to rigorously evaluate the decision-making capabilities of LLMs. Building on the pioneering work of Krishnamurthy et al. (2024), we significantly expand the benchmark by incorporating a broader range of tasks with varying complexities, including variations in the number of arms, reward distributions, exploration difficulty, and textual descriptions of environments. Additionally, we extend it to CB environments, where rewards across arms depend on contextual features, to assess generalization in LLM exploration.
@@
To enhance LLMs for in-context exploration, we leverage known bandit algorithms such as UCB and Thompson Sampling (Thompson, 1933), which have been proven "optimal" under mild conditions. We investigate two approaches: (1) inference-time algorithm-guided support, where summary statistics on interaction history, along with descriptions of bandit algorithms, are provided in context for LLMs to choose actions, and (2) algorithm distillation via optimal demonstration data, where “oracle” trajectories from optimal bandit algorithms are provided as either in-context few-shot demonstrations or oracle behavior fine-tuning. We benchmark off-the-shelf LLMs of different sizes - both open-sourced and proprietary - and those enhanced by our approaches on BanditBench. Both approaches demonstrate promising improvements over baseline methods that rely solely on raw interaction histories presented as sequences of (action, reward) tuples. Furthermore, our results show that fine-tuning to distill optimal exploration behavior leads to strong generalization across domains, enabling smaller models to achieve superior exploration performance compared to larger models. We also perform extensive ablation studies that reveal how training task difficulty, textual representation and Algorithm-Guided Support impact model performance. To gain deeper insights into the exploration efficiency of different methods, we fit a parametric function to the observed regret patterns, allowing for a more rigorous interpretation of the exploration efficiencies of various LLMs and our proposed approaches.
@@ 4.3.2 Results and Ablation studies
Figure 1 presents a comparative overview of in-context few-shot demonstration, oracle behavior fine-tuning, and inference-time algorithmic guidance performance across various model sizes and training configurations. Few-shot demonstrations exhibited contrasting effects on Gemini-1.5 Flash and Pro. While few-shot learning boosts the performance of Flash beyond the best inference-time setup, it surprisingly hurts Pro’s performance in both MAB and CB. Aligned with the observations in Zheng et al. (2024), our hypothesis is that few shot examples we manually crafted could disrupt the CoT structure in these larger models, which requires the few-shot examples to be carefully tuned in order to be helpful. Further analysis reveals the remarkable effectiveness of oracle behavior fine-tuning. It significantly outperforms both few-shot and baseline approaches in both MAB and CB across all model sizes, even larger ones. This robust improvement highlights the effectiveness of directly optimizing model parameters for the exploration task. Notably, the best fine-tuned Gemini-1.5 Flash model surpasses even the highest-performing Gemini-1.5 Pro model. The significant advantage of fine-tuning over few-shot learning and baseline performance highlights its potential as a key technique for enhancing LLM exploration capabilities.
@@
Impact of History Textualization at Inference Time
We examine how different inference-time support techniques—namely RH, SH, and AG—influence model performance, each utilizing distinct history textualization functions $\phi$, as introduced in Section 2. It is worth mentioning that in the MAB setup, both SH and AG significantly reduce context length compared to RH, to $O(K)$ instead of $O(t)$. As illustrated in Table 1, leveraging inference-time support (i.e., SH and AG) significantly enhances exploration performance across all models. This supports the intuition that effective in-context exploration requires more than memorizing input-output pairs; it demands reasoning to extract sufficient statistics from raw data and utilize them effectively for balancing exploration and exploitation. However, the exact benefit of incorporating UCB-style information in the MAB setup remains uncertain. We hypothesize that under MAB, the exploitation value and exploration bonus are straightforward transformations of the empirical mean and the number of times each arm has been pulled $N_{t}(a)$, and that the LLM has the capacity to learn the functional form efficiently. In CB, we compare AG to RH and find a substantial improvement. This gap is particularly significant, as learning the exploitation value and exploration bonus in this scenario requires the model to implicitly solve ridge regression and determine the appropriate functional form of the high-probability confidence bound, making it a more complex reasoning task. The algorithmic guide approach can thus be seen as LLMs calling external tools to compute sufficient statistics required for optimal exploration.
Impact of Task Difficulty in Oracle Trajectories
@@
We examine whether the choice of oracle trajectories used in both in-context demonstration and oracle behavior fine-tuning significantly affects the model’s performance during inference. To investigate this, we select trajectories from two extreme setups. The easiest setup involves (Bernoulli, Video, Large $\Delta_{min}$, $K=5$), denoted as $D_{\text{easy}}$, with AG. Conversely, the hardest setup, denoted as $D_{\text{hard}}$ utilizes (Bernoulli, Clothes, Small $\Delta_{min}$, $K=20$), with RH. Figure 3(a) illustrates that the choice of oracle trajectories significantly impacts the model’s performance, with a surprising contrast between the two algorithm distillation methods. In-context demonstration achieves a higher win-rate when using $D_{\text{easy}}$ as demonstration (50.2) compared to when using $D_{\text{hard}}$ (43.0). This suggests that the limited examples provided in the demonstrations may be insufficient for the model to effectively utilize them under the higher complexity and subtle reward signals of the harder task. Conversely, fine-tuning exhibits the opposite trend, with a higher win-rate when trained on $D_{\text{hard}}$ (65.6) compared to $D_{\text{easy}}$ (54.5). This implies that fine-tuning, with its extensive training data, might be overfitting to the specific nuances of the training distribution, leading to poor generalization when faced with a different task structure.
@@
(b) Textual Representation, RH vs SH (MAB).
(c) Textual Representation with and without AG (CB).
@@
Impact of Textualization in Oracle Trajectories
We further investigate the effect of textualization in the oracle trajectories. We consider two representations in MAB: RH and SH. The results in Figure 3(b) reveal a clear contrast in how these representations affect the two algorithm distillation methods. For in-context demonstration, SH leads to significantly better performance (50.2% win-rate) compared to RH (27.5% win-rate). This suggests that providing concise, informative summaries of optimal exploration behavior is more effective for few-shot learning than presenting the complete raw history. On the other hand, fine-tuning exhibits the opposite trend. RH has a substantially higher win-rate (65.6) compared to SH (28.3). This indicates that fine-tuning benefits from the richer information present in complete action-reward sequences, allowing it to learn more nuanced patterns of the optimal exploration strategy. These contrasting preferences for textual representation in oracle trajectories highlight the nuanced ways in which fine-tuning and few-shot learning interact with different types of information. Furthermore, in CB, we observe a significant impact of incorporating algorithm-guided (AG) information into the oracle trajectories for fine-tuning. Augmenting RH with AG details, including the exploitation value and exploration bonus, leads to a dramatic improvement in win-rate, rising from 28.6 to 89.3 in Figure 3(c). This suggests that providing the LLM with explicit insights into the underlying decision-making process of the oracle algorithm (UCB, in this case), in addition to the complete action-reward sequence, significantly enhances its ability to learn and generalize the optimal exploration strategy in the CB environment.
Figure 4: Impact of Textual Representation at Inference Time.
Impact of Trajectory and Inference-time Representation Alignment
@@ 5 Functional Interpretation of LLM Exploration Behavior
The three parameters $\alpha,\beta,\lambda_{1}$ in the equation are all positive real numbers. $\lambda_{2}$ is unconstrained. $\Delta_{\min}$ captures the gap between the best and second best arm. This functional form provides intuitive interpretations for the underlying parameters. Specifically, $\log(T)$ represents sublinear scaling of the regret, which is known to be achieved by only the best bandit algorithms (e.g. UCB and Thompson Sampling). The $T$ scaling describes a linear growth or the inability of an agent to match the optimal policy $\pi^{*}$. This means a strong algorithm should have $\alpha$ as small as possible, and have $\beta=0$. This functional form also allows us to see some growth behaviors in-between with both positive $\alpha$ and $\beta$. We use the curve fit function in Scikit-learn (Pedregosa et al., 2011) to fit the cumulative regret curve of UCB and LLMs coupled with different methods (i.e., inference-time algorithm-guided support, in-context demonstration, and oracle behavior fine-tuning). The results of the fitted $\alpha$ and $\beta$ values are presented in Figure 5. For the largest Pro models, applying effective inference-time support, such as AG and SH can achieve nearly sub-linear regret. More intriguingly, for Flash models, fine-tuning for oracle behavior significantly boosts performance, enabling them to attain sub-linear regret with a lower $\alpha$. In contrast, weaker models such as Gemma 2B and 9B appear to remain in the linear regret regime across nearly all methods.

### v2

@@ Abstract
Despite their success in many domains, large language models (LLMs) remain under-studied in scenarios requiring optimal decision-making under uncertainty. This is crucial as many real-world applications, ranging from personalized recommendations to healthcare interventions, demand that LLMs not only predict but also actively learn to make optimal decisions through exploration. In this work, we measure LLMs’ (in)ability to make optimal decisions in bandits, a state-less reinforcement learning setting relevant to many applications. We develop a comprehensive suite of environments, including both context-free and contextual bandits with varying task difficulties, to benchmark LLMs’ performance. Motivated by the existence of optimal exploration algorithms, we propose efficient ways to integrate this algorithmic knowledge into LLMs: by providing explicit algorithm-guided support during inference; and through algorithm distillation via in-context demonstrations and fine-tuning, using synthetic data generated from these algorithms. Impressively, these techniques allow us to achieve superior exploration performance with smaller models, surpassing larger models on various tasks. We conducted an extensive ablation study to shed light on various factors, such as task difficulty and data representation, that influence the efficiency of LLM exploration. Additionally, we conduct a rigorous analysis of the LLM’s exploration efficiency using the concept of regret, linking its ability to explore to the model size and underlying algorithm. The code is available at https://github.com/allenanie/EVOLvE.
@@ 1 Introduction
We study LLMs’ in-context exploration capabilities under the simplified framework of bandits — a stateless form of reinforcement learning that is highly applicable to many domains. We set up the LLM to interact with the environment over $T$ rounds. In each round, it receives the full history of its past interactions, the current state (if provided), and a set of actions, and it is tasked with selecting an action to maximize the cumulative reward. Ideally, the LLM should adaptively choose an action in each round to learn the reward distributions of different actions and eventually converge to consistently selecting the optimal one. We study LLM’s ability to do so in-context, without the need to re-train, which we dubbed as in-context exploration. Unlike using LLM for reward-free or curiosity-driven exploration, in-context exploration is in-context self-improvement, where the LLM builds up context through interactions to solve a task.
We introduce BanditBench [^1] Install the library: pip install banditbench, a comprehensive suite of multi-armed bandit (MAB) (Slivkins et al., 2019) and contextual bandit (CB) (Li et al., 2010) environments in natural language to rigorously evaluate the decision-making capabilities of LLMs. Building on the pioneering work of Krishnamurthy et al. (2024), we significantly expand the benchmark by incorporating a broader range of tasks with varying complexities, including variations in the number of arms, reward distributions, exploration difficulty, and textual descriptions of environments. Additionally, we extend it to CB environments, where rewards across arms depend on contextual features, to assess generalization in LLM exploration.
@@
To enhance LLMs for in-context exploration, we leverage known bandit algorithms such as Upper-Confidence Bound (UCB) algorithm, which have been proven "optimal" under mild conditions. We investigate two approaches: (1) inference-time algorithm-guided support, where summary statistics on interaction history, along with descriptions of bandit algorithms, are provided in context for LLMs to choose actions, and (2) algorithm distillation via optimal demonstration data, where “oracle” trajectories from optimal bandit algorithms are provided as either in-context few-shot demonstrations or optimal behavior fine-tuning. We benchmark off-the-shelf LLMs of different sizes - both open-sourced and proprietary - and those enhanced by our approaches on BanditBench. Both approaches demonstrate promising improvements over baseline methods that rely solely on raw interaction histories presented as sequences of (action, reward) tuples. Furthermore, our results show that fine-tuning to distill optimal exploration behavior leads to strong generalization across domains, enabling smaller models to achieve superior exploration performance compared to larger models. We also perform extensive ablation studies that reveal how training task difficulty, textual representation and Algorithm-Guided Support impact model performance. To gain deeper insights into the exploration efficiency of different methods, we fit a parametric function to the observed regret patterns, allowing for a more rigorous interpretation of the exploration efficiencies of various LLMs and our proposed approaches.
@@ 4.2 Algorithm Distillation via Demonstration and Fine-tuning
Optimal Behavior Fine-tuning (OFT) and Behavior Cloning (BC) (Pomerleau, 1991) share many similarities. However, they are not the same. Although both approaches rely on maximum-likelihood learning, their objectives are different: OFT seeks to encode a dynamic, iterative refinement process, while BC focuses on replicating static behavior. OFT is designed for algorithm distillation, focusing on capturing a sequence of self-improvement behaviors, and generalization across any new test domains. In contrast, BC aims to learn a policy by mimicking a static policy, with no iterative improvement between trajectories.
This difference becomes very clear when we work through this example. Given a policy $\pi$, if our goal is to learn a policy $\hat{\pi}$ that perfectly mimics the behavior of $\pi$, then we will create a behavior cloning dataset ${\mathcal{D}}_{\text{BC}}$, by collecting this policy $\pi$’s interaction with the environment. In this dataset, for the same state $s$, the policy remains unchanged, which the means $\pi(a|s)$ remains the same in the entire dataset given the same state $s$. However, if our goal is to learn to mimic the improvement of this policy $\pi$ as it interacts with the environment – note that the policy improves after a policy improvement operation (e.g., policy gradient update, or q-value update), we are constructing a dataset ${\mathcal{D}}_{\text{OFT}}$ from a set of policies: $\{\pi_{1},\pi_{2},...,\pi_{*}\}$. This means, given the same state $s$, the action to take will change: $\pi_{i}(a|s)$, depending on whether this is an early or a late trajectory of this dataset. In order to encode the policy change $i$, the prediction target is $a$ given the current state $s$ and all the previous experiences that were used to improve this policy.
@@ 4.3.2 Results and Ablation studies
Figure 1 presents a comparative overview of in-context few-shot demonstration, oracle behavior fine-tuning, and inference-time algorithmic guidance performance across various model sizes and training configurations. Detailed numbers are reported in Table 1 and 2. Few-shot demonstrations exhibited contrasting effects on Gemini-1.5 Flash and Pro. While few-shot learning boosts the performance of Flash beyond the best history contextualization setup, it hurts Pro’s performance in both MAB and CB. Aligned with the observations in Zheng et al. (2024); Guo et al. (2025), our hypothesis is that few-shot examples we manually crafted could disrupt the CoT structure in larger models, which requires the few-shot examples to be carefully tuned in order to be helpful. Further analysis reveals the remarkable effectiveness of oracle behavior fine-tuning. It significantly outperforms both few-shot and baseline approaches in both MAB and CB across all model sizes, even larger ones. This robust improvement highlights the effectiveness of directly optimizing model parameters for the exploration task. Notably, the best fine-tuned Gemini-1.5 Flash model surpasses even the Gemini-1.5 Pro model, highlighting its potential as a key technique for enhancing LLM exploration capabilities.
@@
Flash (RH) Flash (AG) Pro (RH) Pro (AG)
MAB, K=5 33.6 26.6 48.0 67.6
MAB, K=20 21.9 37.9 43.0 51.6
CB, K=10 0.0 35.7 7.1 57.1
CB, K=30 0.0 57.1 7.1 71.4
Table 3: Situations where Algorithm-Guided Support significantly outperforms Raw History during inference.
Algorithm-Guided Support Helps on Complex Environments
We observe consistent improvements when transitioning from RH to AG in two scenarios: (1) larger models like Flash and Pro, and (2) more complex scenarios, such as contextual bandits. Table 3 shows that AG provided consistent help when the number of actions is large for both Flash and Pro models. We hypothesize that providing AG is crucial when the action space is large or when decision scenarios are complex.
Impact of Task Difficulty
@@
We examine whether the choice of oracle trajectories used in both few-shot demonstration and oracle behavior fine-tuning affects the model’s performance during inference. To investigate this, we select trajectories from two extreme setups. The easiest setup involves (Bernoulli, Video, Large $\Delta_{min}$, $K=5$), denoted as $D_{\text{easy}}$, with AG. Conversely, the hardest setup, denoted as $D_{\text{hard}}$ utilizes (Bernoulli, Clothes, Small $\Delta_{min}$, $K=20$), with RH. Figure 3(a) illustrates that the choice of oracle trajectories significantly impacts the model’s performance, with a surprising contrast between the two algorithm distillation methods. Few-shot demonstration achieves a higher win-rate when using $D_{\text{easy}}$ as demonstration (50.2) compared to when using $D_{\text{hard}}$ (43.0). This suggests that the limited examples provided in the demonstrations may be insufficient for the model to effectively utilize them under the higher complexity and subtle reward signals of the harder task. Conversely, fine-tuning exhibits the opposite trend, with a higher win-rate when trained on $D_{\text{hard}}$ (65.6) compared to $D_{\text{easy}}$ (54.5). This implies that fine-tuning, with its extensive training data, might be overfitting to the specific nuances of the training distribution, leading to poor generalization when faced with a different task structure.
Impact of Contextualization
We further investigate the effect of contextualization in oracle trajectories. We consider two representations in MAB: RH and SH. Results in Figure 3(b) reveal a clear contrast in how these representations affect the two algorithm distillation methods. For few-shot demonstration, SH leads to significantly better performance (50.2% win-rate) compared to RH (27.5% win-rate). This suggests that providing concise, informative summaries of optimal exploration behavior is more effective for few-shot learning than presenting the complete raw history. On the other hand, fine-tuning exhibits the opposite trend. RH has a substantially higher win-rate (65.6) compared to SH (28.3). This indicates that fine-tuning benefits from the richer information present in complete action-reward sequences, allowing it to learn more nuanced patterns of the optimal exploration strategy. These contrasting preferences for textual representation in oracle trajectories highlight the nuanced ways in which fine-tuning and few-shot learning interact with different types of information. Furthermore, in CB, we observe a significant impact of incorporating AG information into the oracle trajectories for fine-tuning, leading a dramatic improvement in win-rate, rising from 28.6 to 89.3 (Figure 3(c)). This suggests that providing LLMs with explicit insights, in addition to the complete action-reward sequence, enhances its ability to learn and generalize the optimal exploration strategy in the CB environment.
@@
(b) History Context Representation, RH vs SH (MAB).
(c) History Context Representation with and without AG (CB).
@@
Algorithm Distillation Generalizes
We also conducted experiments to evaluate the distillation algorithm’s ability to generalize from smaller to larger action spaces (i.e., easy-to-hard domain generalization). This analysis offers insights into the scalability and adaptability of our approach to more complex domains, such as real-world recommendation systems and other complex decision-making problems. We report this in Table 4. Using few-shot demonstrations or doing oracle behavior fine-tuning from simple-domain trajectories collected (Bernoulli, Easy, K=5) can indeed learn exploration strategies that generalize to a harder domain (Bernoulli, Hard, K=20).
MAB
Hard
Flash
Flash + Few-Shot (SH)
($\mathcal{D}_{\text{easy}}$ on K=5)
Flash + OFT (RH)
($\mathcal{D}_{\text{easy}}$ on K=5)
K=20 34.0% 43.0% 48.4%
Table 4: Easy-to-Hard Generalization: We collect distillation data from easy bandit tasks to learn the exploration behavior and evaluate on hard bandit tasks. History representation during evaluation is in parentheses.
(a) In-Context Few-Shot Demonstration
(b) Oracle Behavior Fine-Tuning
Figure 4: The Impact of Context Representation in In-Context Examples and Fine-Tuning Data at Inference Time. We vary the context representation during “learning” (in-context and weight fine-tuning) and provide or withhold information during inference. Providing AG during inference will always increase the model’s performance. Providing it during learning allows the model to maximally leverage such infomation.
Impact of Training and Inference-time Representation Alignment
@@
Optimality Analysis of Exploration
We use two metrics to capture high-level exploration behaviors: (1) MinFrac, proposed by Krishnamurthy et al. (2024), captures the fraction of pulls allocated to the least-selected arm. An ideal algorithm should exhibit high MinFrac during early exploration and gets lower as $t$ increases, indicating effective exploration; (2) OptFrac, tracks the percentage of times the optimal arm is pulled. Ideally, this percentage should increase as the process progresses, indicating the model’s ability to self-improve. Using Flash model and Bernoulli bandit configurations as an example in Table 5, we see that with AG, the Flash model on average explores more compared to SH and have higher success at identifying the best arm.
MinFrac / $t$ Step 100-th 250-th 500-th 750-th 1000-th
UCB 82.3% 48.6% 27.8% 19.6% 15.3%
Flash (SH) 10.2% 4.2% 2.1% 1.4% 1.1%
Flash (AG) 11.3% 4.5% 2.3% 1.5% 1.1%
OptFrac
UCB 32.7% 49.4% 58.7% 62.6% 65.0%
Flash (SH) 9.3% 10.1% 10.4% 10.6% 10.7%
Flash (AG) 14.4% 15.6% 16.3% 16.6% 16.8%
Table 5: Optimality of Exploration: We show the measures on the $t$-th step along the progression of exploration, averaged across Bernoulli MAB configurations. Full version with more models in Table A4 and A5.
While a model might achieve high performance by chance—e.g., consistently selecting the optimal arm without deliberate exploration—we examine its behavior more closely using two metrics. OptFrac measures how often the optimal arm is chosen, revealing whether the model increasingly favors the best option. As shown in Table 5, UCB’s OptFrac steadily rises over time (32.7% $\rightarrow$ 65.0%), while Gemini-1.5 Flash remains nearly flat (9.3% $\rightarrow$ 10.7%), suggesting limited adaptation. To assess directed exploration, we use MinFrac, the fraction of pulls allocated to the least-selected arm. UCB shows a desirable decline (82.3% $\rightarrow$ 15.3%), indicating early broad exploration followed by exploitation. In contrast, Gemini-1.5 Flash starts low and drops quickly (11.3% $\rightarrow$ 1.1%), implying minimal directed exploration throughout. Together, these metrics demonstrate that Gemini’s performance is not driven by meaningful exploration. More discussion in Appendix A.6.
@@ 5 Functional Interpretation of LLM Exploration Behavior
The three parameters $\alpha,\beta,\lambda_{1}$ in the equation are all positive real numbers. $\lambda_{2}$ is unconstrained. $\Delta_{\min}$ captures the gap between the best and second best arm. This functional form provides intuitive interpretations for the underlying parameters. Specifically, $\log(T)$ represents sublinear scaling of the regret, which is known to be achieved by only the best bandit algorithms (e.g. UCB and Thompson Sampling). The $T$ scaling describes a linear growth or the inability of an agent to match the optimal policy $\pi^{*}$. This means a strong algorithm should have $\alpha$ as small as possible, and have $\beta=0$. This functional form also allows us to see some growth behaviors in-between with both positive $\alpha$ and $\beta$. We use the curve fit function in Scikit-learn (Pedregosa et al., 2011) to fit the cumulative regret curve of UCB and LLMs coupled with different methods (i.e., inference-time algorithm-guided support, in-context demonstration, and optimal behavior fine-tuning). The results of the fitted $\alpha$ and $\beta$ values are presented in Figure 5. For the largest Pro models, applying effective inference-time support, such as AG and SH can achieve nearly sub-linear regret. More intriguingly, for Flash models, fine-tuning for optimal behavior significantly boosts performance, enabling them to attain sub-linear regret with a lower $\alpha$. In contrast, weaker models such as Gemma 2B and 9B appear to remain in the linear regret regime across nearly all methods.